OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. (9,-7); 7x-4y=5 the equation of the line is y=?

OpenStudy (campbell_st):

for the line 7x -4y = 5 make y the subject so you can identify the slope of the line...

OpenStudy (phi):

add +4y to both sides as a first step 7x -4y + 4y = 4y+5 can you simplify the left side (of the = sign) ?

OpenStudy (anonymous):

ok so how do I do thst then? im so confused

OpenStudy (phi):

-4 dollars plus 4 dollars -4 cows plus 4 cows or 4 cents - 4 cents 4 y's - 4 y's you get the same answer for each of those .

OpenStudy (anonymous):

okay so how wouls I solve it to find it

OpenStudy (anonymous):

equation of any line parallel to 7x-4y=5 is 7x-4y=c because it passes through (9,-7) put in the above eq.and find the value of c and then finally put the value of c.